OpenStudy (anonymous):

My brother is 4 times as old as I am. Six years from now, he will be twice as old as I am. How old is my brother now?

OpenStudy (anonymous):

Brother = 4(Me) Brother + 6 = 2(Me+6) 4(Me) + 6 = 2(Me) + 12 2(Me) = 6 (Me) = 3 Brother = 12